# LLaVA-ReID: Selective Multi-image Questioner for Interactive Person Re-Identification

### v2

@@ 1 Introduction
  • We introduce a new task, Interactive Person Re-Identification, which goes beyond traditional T-ReID by leveraging interaction with witnesses to improve accuracy and adaptability. To support this task, we construct a tailored dataset featuring multi-round dialogues, enabling more effective training and evaluation of Inter-ReID systems.
@@ 2.1 Text-to-Image Person Re-identification
Person re-identification has gained significant attention as a critical task in surveillance and security systems (Ye et al., 2021), focusing on matching individuals across non-overlapping camera views (Yang et al., 2022) and varying modalities (Ye et al., 2020; Gong et al., 2024; Shi et al., 2024). Among the diverse Re-ID challenges, text-to-image ReID aims to retrieve the corresponding image of a person based on textual descriptions. The primary challenge lies in effectively aligning fine-grained visual features with language descriptions in a joint embedding space. Existing methods can be broadly categorized into two approaches: network architecture design and training objective design. The first category  (Li et al., 2017; Gao et al., 2021; Shao et al., 2022) focuses on developing fine-grained fusion networks such as multi-scale interaction (Yan et al., 2023) mechanisms to encourage the alignment of visual details. The second category introduces auxiliary tasks to improve the fine-grained alignment between body regions and textual entities, such as random masking (Jiang & Ye, 2023; Bai et al., 2023), color prediction (Wu et al., 2021), and attribute prediction (Zuo et al., 2023).
These approaches typically assume that descriptions are complete and well-structured. However, in real-world scenarios, witnesses rarely provide a detailed and comprehensive account of the target person’s appearance, often resulting in partial and vague descriptions. To address this limitation, we propose an interactive ReID framework that incorporates multiple rounds of dialogue between the witness and the retrieval system. This iterative process enables the system to refine the initial description by actively engaging the witness with targeted questions. Through active questioning, our system uncovers missing details thus improving retrieval performance. While some recent works (Tan et al., 2024; Zuo et al., 2024, 2023) attempt to annotate more fine-grained descriptions, they focus on improving fine-grained alignment rather than guiding the witness to supplement overlooked details through a dialogue-based approach.
@@ 2.2 Interactive Cross-Modal Retrieval
Interactive cross-modal retrieval involves refining the initial query by engaging the user to provide additional information or clarifications. Several studies have introduced interactive mechanisms in text-to-image retrieval (Levy et al., 2023; Lee et al., 2024; Zhu et al., 2024) and text-to-video retrieval (Madasu et al., 2022; Liang & Albanie, 2023), investigating various interaction formats (Kovashka et al., 2015; Cai et al., 2021; Lee et al., 2021). Among these formats, question-answering using free-form text dialogue has shown the most promise, as it closely mirrors the natural way humans interact (Madasu et al., 2022). For example, SimIRV (Liang & Albanie, 2023) designs a heuristic question generator that asks about the attributes of objects in the image. Inspired by the context-aware capabilities of large language models (LLMs), ChatIR (Levy et al., 2023) leverages LLMs to generate the next question based on few-shot examples. More recently, PlugIR (Lee et al., 2024) proposes further enhancing performance by filtering redundant questions and rewriting the user query.
However, these methods primarily focus on general application scenarios and lack the domain-specific knowledge required for fine-grained retrieval tasks. To address this, we introduce the first customized dataset for training and evaluating Inter-ReID. The dataset includes subtle clothing features, accessories, and other distinctive attributes, enabling the questioner to effectively handle the nuances between individuals. Furthermore, we propose a question generator based on the large multi-modal model (Li et al., 2024b), which integrates both visual and textual contexts from candidate images and dialogue inputs. In contrast, existing methods first transform images into textual descriptions using image caption models (Li et al., 2022), which might lack the fine-grained visual details and multi-modal perception required for context-aware question generation.
@@ 3.2 Dataset Construction
This step divides fine-grained descriptions into sub-captions to facilitate the construction of question-answer pairs in Step 3. Two key principles guide the creation of sub-captions: i) they must address specific details not mentioned in the initial description, and ii) each sub-caption should focus on a single and distinct aspect of the person. To achieve this, we first remove the information already provided in the initial description to generate a follow-up description that emphasizes the unexplored details. Next, we decompose the follow-up description into non-overlapping sub-captions using the prompt in Table 7. Each sub-caption represents a unique aspect of the person’s features and ensures that no redundant questions would be asked.
@@ 5.3 Integration with Text-based ReID
Table 3: Analysis of image and question selection.
@@
Candidate k-means 72.51 90.08 94.59 1.049
uniform 72.70 90.00 94.60 1.060
top-k 72.53 89.78 94.22 1.079
@@ 6 Conclusion
This paper introduces Interactive Person Re-Identification, a new problem that more closely aligns with real-world scenarios by incorporating multi-round dialogue between the witness and the retrieval system. To support this task, we propose a specific dataset Interactive-PEDES and propose a multi-image questioner LLaVA-ReID that leverages both visual and textual contexts to identify fine-grained differences. Despite achieving strong performance, challenges remain, such as the limited ability of retriever to handle incremental and dialogue-based descriptions. Future research could further explore these directions.

### v3

@@ 1 Introduction
  • We introduce a new task, Interactive Person Re-Identification, which goes beyond traditional T-ReID by incorporating interaction with witnesses to improve both accuracy and adaptability. To support this task, we construct a tailored dataset featuring multi-round dialogues, enabling more effective training and evaluation of Inter-ReID systems.
@@ 2.1 Text-to-Image Person Re-identification
Person re-identification has gained significant attention as a critical task in surveillance and security systems (Ye et al., 2021), focusing on matching individuals across non-overlapping camera views (Yang et al., 2022) and varying modalities (Ye et al., 2020; Gong et al., 2024; Shi et al., 2024). Among the diverse Re-ID challenges, text-to-image ReID aims to retrieve the corresponding image of a person based on textual descriptions. The primary challenge lies in effectively aligning fine-grained visual features with language descriptions in a joint embedding space. Existing methods can be broadly categorized into two approaches: network architecture design and training objective design. The first category  (Li et al., 2017; Gao et al., 2021; Shao et al., 2022) focuses on developing fine-grained fusion networks such as multi-scale interaction (Yan et al., 2023) mechanisms to encourage the alignment of visual details. The second category introduces auxiliary tasks to improve the fine-grained alignment between body regions and textual entities, such as random masking (Jiang & Ye, 2023; Bai et al., 2023), color prediction (Wu et al., 2021; Gong et al., 2022), and attribute prediction (Zuo et al., 2023).
These approaches typically assume that descriptions are complete and well-structured. However, in real-world scenarios, witnesses rarely provide a detailed and comprehensive account of the target person’s appearance, often resulting in partial and vague descriptions. To address this limitation, we propose an interactive ReID framework that incorporates multiple rounds of dialogue between the witness and the retrieval system. This iterative process enables the system to refine the initial description by actively engaging the witness with targeted questions. Through active questioning, our system uncovers missing details, thus improving retrieval performance. While some recent works (Tan et al., 2024; Zuo et al., 2024, 2023) attempt to annotate more fine-grained descriptions, they focus on improving fine-grained alignment rather than guiding the witness to supplement overlooked details through a dialogue-based approach.
@@ 2.2 Interactive Cross-Modal Retrieval
Interactive cross-modal retrieval involves refining the initial query by engaging the user to provide additional information or clarification. Several studies have introduced interactive mechanisms in text-to-image retrieval (Levy et al., 2023; Lee et al., 2024; Zhu et al., 2024) and text-to-video retrieval (Madasu et al., 2022; Liang & Albanie, 2023), investigating various interaction formats (Kovashka et al., 2015; Cai et al., 2021; Lee et al., 2021; Ding et al., 2025). Among these formats, question-answering using free-form text dialogue has shown the most promise, as it closely mirrors the natural way humans interact (Madasu et al., 2022). For example, SimIRV (Liang & Albanie, 2023) designs a heuristic question generator that asks about the attributes of objects in the image. Inspired by the context-aware capabilities of large language models (LLMs), ChatIR (Levy et al., 2023) leverages LLMs to generate the next question based on few-shot examples. More recently, PlugIR (Lee et al., 2024) proposes further enhancing performance by filtering redundant questions and rewriting the user query.
However, these methods primarily focus on general application scenarios and lack the domain-specific knowledge required for fine-grained retrieval tasks (Lin et al., 2023). To address this, we introduce the first customized dataset for training and evaluating Inter-ReID. The dataset includes subtle clothing features, accessories, and other distinctive attributes, enabling the questioner to effectively handle the nuances between individuals. Furthermore, we propose a question generator based on the large multi-modal model (Li et al., 2024b), which integrates both visual and textual contexts from candidate images and dialogue inputs. In contrast, existing methods first transform images into textual descriptions using image caption models (Li et al., 2022), which might lack the fine-grained visual details and multi-modal perception required for context-aware question generation.
@@ 3.2 Dataset Construction
This step divides fine-grained descriptions into sub-captions to facilitate the construction of question-answer pairs in Step 3. Two key principles guide the creation of sub-captions: i) they must address specific details not mentioned in the initial description, and ii) each sub-caption should focus on a single and distinct aspect of the person. To achieve this, we first remove the information already provided in the initial description to generate a follow-up description that emphasizes the unexplored details. Next, we decompose the follow-up description into non-overlapping sub-captions using the prompt in Table 7. Each sub-caption represents a unique aspect of the person’s features and ensures that no redundant questions will be asked.
@@ 5.3 Integration with Text-based ReID
Table 3: Analysis of image and question selection strategy.
@@
Candidate k-means 72.51 90.08 94.59 0.727
uniform 72.70 90.06 94.60 0.735
top-k 72.53 89.78 94.22 0.740
@@ 6 Conclusion
This paper introduces Interactive Person Re-Identification, a new problem that more closely aligns with real-world scenarios by incorporating multi-round dialogue between the witness and the retrieval system. To support this task, we propose a specific dataset Interactive-PEDES and propose a multi-image questioner LLaVA-ReID that leverages both visual and textual contexts to identify fine-grained differences. Despite achieving strong performance, challenges remain, such as the limited ability of the retriever to handle incremental and dialogue-based descriptions. Future research could further explore these directions.
@@
Acknowledgments
This work was supported in part by NSFC under Grant 62176171, 624B2099, 62472295, U24B20174; in part by the Fundamental Research Funds for the Central Universities under Grant CJ202303, CJ202403; in part by Sichuan Science and Technology Planning Project under Grant 24NSFTD0130; and in part by Baidu Scholarship.